# SPECIAL RELATIVITY

### v11

@@ 16.1 Manifestly Covariant Formulation
The covariant equation of motion for a relativistic charged particle, subject to the four-force $K_{\mu}=eF^{\mu\nu}dx_{\nu}/d\tau$, as given in Eq. (21), represents a relativistic extension of Newton’s second law. While the classical form $md\vec{v}/dt=\vec{f}$ remain valid in the particle’s instantaneous comoving Lorentz frame, the relativistic formulation embeds these three spatial equations into the four-dimensional Minkowski spacetime  [^52] The notion of embedding relies on the fact that Newton’s second law can always be applied in any Lorentz frame in which the particle is instantaneously at rest [64]. That is, if a comoving Lorentz frame is defined at a given instant, the particle’s evolution can be accurately predicted in this frame over an infinitesimal time interval. Geometrically, Newton’s law holds rigorously on a hyperplane orthogonal to the particle’s world line. As the particle moves, the hyperplane—and its normal vector $u_{\mu}$—tilts accordingly along the world line. To formalize this embedding, one introduces a projection operator $\hat{P}_{\perp}$ that continuously projects vectors in Minkowski space onto hyperplanes orthogonal to the world line. This operator is given by $(\hat{P}_{\perp})_{\mu\nu}=\eta_{\mu\nu}-u_{\mu}u_{\nu}/u^{2}$ [64]. The presence of $\hat{P}_{\perp}$ highlights that only three of the four components of the covariant equation are independent. .
@@ 16.2 Conventional Particle Tracking
It should be noted that usually the notation ”$\tau$” arose from the proper time. Now let us recall the standard concept of the object’s proper time. Let a point-like object move uniformly and rectilinearly relative to the lab frame $K$. The proper frame $K^{\prime}$ can be fixed to the moving object. The object is at rest in this frame, so that events happening with this object are registered by one clock. This clock counts the proper time at the point where the object is located. As we already know, the proper time on moving object flows slower than the time $t$. This phenomenon was called time dilation. The proper time can also be introduced for a point particle moving with acceleration. This standard interpretation of the proper time $\tau$ has nothing to do with dynamics and one may wonder where this contradiction existing between the name and the content of time in dynamic law comes from.
@@ 16.4 Standard Integration of the 4D Covariant Equation of Motion
We begin by analyzing the textbook’s treatment of the integration of the covariant equation of motion. Consider, for example, the motion of a charged particle in a given electromagnetic field. Of particular practical importance is the simplest case: a uniform electromagnetic field, meaning that the field tensor $F^{\mu\nu}$ remains constant throughout the entire space-time region of interest. Specifically, we focus on the motion of a particle in a constant, homogeneous magnetic field, represented by the tensor components $F^{\mu\nu}=B(e^{\mu}_{2}e^{\nu}_{3}-e^{\nu}_{2}e^{\mu}_{3})$ where $e^{\mu}_{2}$ and $e^{\mu}_{3}$ are orthonormal space like basis vectors $e^{2}_{2}=e^{2}_{3}=-1$, $e_{2}\cdot e_{3}=0$. In the lab frame of reference where $e^{\mu}_{0}$ is taken as the time axis, and $e^{\mu}_{2}$ and $e^{\mu}_{3}$ are space vectors the field is indeed purely magnetic, of magnitude $B$ and parallel to the $e_{1}$ axis. Let us set the initial four-velocity $u^{\mu}(0)=\gamma ce^{\mu}_{0}+\gamma ve^{\mu}_{2}$, where $v$ is the initial particle’s velocity relative to the lab observer along the axis $e_{2}$ at the instant $\tau=0$, and $\gamma=1/\sqrt{1-v^{2}/c^{2}}$. The components of the equation of motion are then $du^{(0)}/d\tau=0$, $du^{(1)}/d\tau=0$, $du^{(2)}/d\tau=-eBu^{(3)}/(mc)$, $du^{(3)}/d\tau=eBu^{(2)}/(mc)$. We seek the initial value solution to these equations as done in the existing literature  [^53] Textbook authors, such as those in [5, 64, 65], conclude from the covariant equations that a charged particle in a constant magnetic field moves in a uniform circular trajectory. This analysis does not account for relativistic kinematics, and instead, the Galilean vector law of velocity addition is applied. The non-relativistic kinematics are derived from the relation $d\tau=dt/\gamma$. It is only after making this substitution for $d\tau$ hat the authors obtain the familiar formula for the non-covariant trajectory.. A distinctive aspect of relativistic dynamics is the presence of constraints. The evolution of the particle is subject to $mdu^{\mu}/d\tau=eF^{\mu\nu}u_{\nu}$, but also to the constraint $u^{2}=c^{2}$. However, such a condition can be weakened requiring its validity at certain values of $\tau$ only, let us say initially, at $\tau=0$. Therefore, if $Y(\tau)=u^{2}-c^{2}$ vanishes initially, i.e. $Y(0)=0$, then $Y(\tau)=0$ at any $\tau$. In other words, the differential Lorentz-force equation implies the constraint $u^{2}=c^{2}$ once this is satisfied initially. Integrating with respect to the proper time we have $u^{\mu}(\tau)=\gamma e^{\mu}_{0}+\gamma v[e^{\mu}_{2}\cos(\omega\tau)+e^{\mu}_{3}\sin(\omega\tau)]$ where $\omega=eB/(mc)$. We see that $\gamma$ is constant with time, meaning that the energy of a charged particle moving in a constant magnetic field is constant. After two successive integrations we have $X^{\mu}(\tau)=X^{\mu}(0)+\gamma c\tau e^{\mu}_{0}+R[e^{\mu}_{2}\sin(\omega\tau)-e^{\mu}_{3}\cos(\omega\tau)]$ where $R=\gamma v/\omega$. This enables us to find the time dependence $[0,X^{(2)}(t),X^{(3)}(t)]$ of the particle’s position since $t/\gamma=\tau$. From this solution, we conclude that the motion of a charged particle in a constant magnetic field is indeed uniform circular motion, consistent with classical expectations, yet derived within the fully covariant relativistic framework.
@@
One might expect that the particle’s trajectory in the lab frame, derived from the previous reasoning $[0,X^{(2)}(t),X^{(3)}(t)]$, should be identified with $\vec{x}_{cov}(t)$. However, the trajectory $[0,X^{(2)}(t),X^{(3)}(t)]$ does not include relativistic kinematics effects. We have found that the standard integration of the four-dimensional covariant equation of motion, as shown in Eq. (21), yields a particle trajectory that is mathematically identical to the one derived from Newtonian kinematics and dynamics. Specifically, the trajectory of the electron fails to incorporate relativistic effects, and instead, the Galilean law of velocity addition is applied.
@@ 16.5 Convention-Invariant Particle Tracking
16.6 Lorentz-Covariant Particle Tracking
@@
One of the consequences of the non-commutativity of non-collinear Lorentz boosts is the absence of a shared, common notion of ordinary space. Suppose that in the Lorentz frame $K^{\prime}$, a traveler observes a particle moving along the $y^{\prime}$ axis. In other words, the particle has a velocity component $v^{\prime}_{y}$ in this frame, while the frame $K^{\prime}$ itself is moving with velocity $V$ along the $x$ axis relative to another Lorentz frame $K$. To an observer at rest in frame $K$, the moving frame $K^{\prime}$ appears rotated with respect to $K$. Let us consider the simple case where the boost velocity $V$ approaches the speed of light (i.e., the ultrarelativistic limit), while the transverse velocity $v^{\prime}_{y}$ remains small. This means we are taking the limit where $\gamma\gg 1\gg v^{\prime}_{y}/c$, treating $v^{\prime}_{y}/c$ as a small parameter and keeping terms up to second order. In this regime, the axes of the $K^{\prime}$ frame appear rotated with respect to those of $K$ by an angle $\theta_{w}=v^{\prime}_{y}/(\gamma c)$. Notably, this rotation is in the same direction as the rotation of the particle’s velocity vector in $K$  [^54] To determine the rotation angle in the given problem, we consider a composition of Lorentz boosts. Two successive boosts—first from frame $K$ to $K^{\prime}$, and then an infinitesimal boost to the particle’s proper frame—are equivalent to a single boost from $K$ directly to the proper frame, followed by a rotation. The proper frame is rotated by an angle $d\theta_{w}=dv^{\prime}_{y}/(\gamma c)$ relative to frame $K$. The coordinate axes of the proper frame remain parallel to those of frame $K^{\prime}$. From the perspective of frame $K$, the coordinate axes of frame $K^{\prime}$ are also rotated by the angle $d\theta_{w}$ with respect to those of frame $K$. We now consider the physical consequences of this rotation. First, note that the transverse velocity component $v^{\prime}_{y}$ acquires a projection onto the $x$-axis when viewed from frame $K$. As a result, the component of the particle’s velocity along the $x$-axis in frame $K$ is slightly reduced from $V$.
@@
This result stands in contrast with the textbook prediction, which states that the total speed of the particle in frame $K$ increases from $V$ to $V+v_{y}^{2}/(2c)$. Our result contradicts this claim. The standard derivation of velocity addition relies on the relation $dt^{\prime}=dt/\gamma$, but neglects the mixing of transverse spatial coordinates with time in Lorentz transformations. This omission leads to an incorrect conclusion about the magnitude of the resulting velocity.
16.7 Relativistic Velocity Composition and the Wigner Rotation
Solving problems related to covariant particle tracking presents two major challenges.
The first challenge arises from the conventional approach to particle tracking within a single inertial frame, which is based entirely on classical Newtonian mechanics, without incorporating Lorentz transformations. Standard textbooks assert that when only one frame is involved, relativity is unnecessary. In this framework, the only relativistic modification is the introduction of relativistic mass, while space and time remain distinct, with no mixing of spatial and temporal components. Surprisingly, this crucial aspect has received little attention within the physics community.
The second challenge stems from confusion surrounding the composition of non-parallel velocities in special relativity. In Newtonian mechanics, velocity addition follows the parallelogram rule, treating velocities as ordinary vectors. However, in relativistic kinematics, velocity composition must be derived from the combination of Lorentz boosts. A well-known issue is that Lorentz boosts alone do not form a group, which raises a fundamental concern: if a set of transformations lacks group structure, the resulting theory is mathematically incomplete.
@@ 16.7 Relativistic Velocity Composition and the Wigner Rotation
Textbook authors often assume that the principle of relativity demands that Einstein’s velocity addition law obeys group composition. However, the non-associativity of non-collinear velocity addition is not widely recognized. Many physicists, having learned relativity from standard textbooks, might find this claim surprising, as Einstein’s velocity addition is often assumed to be associative. Yet a straightforward algebraic exercise demonstrates its non-associativity. As a result, constructing a covariant trajectory, $\vec{x}_{cov}(t)$, using only Lorentz boosts is impossible—additional elements are required.
@@
The root cause of this issue lies in the presence of Wigner rotation. In fact, the breakdown of commutativity and associativity in Einstein’s velocity addition law naturally manifests as Wigner rotation, offering deeper insight into the structure of relativistic transformations. The next section will explore the relativistic transformation of velocities in greater detail.
16.8 Mathematical Analysis of Relativistic Velocity Composition
Above, we derived specific quantitative results for the relativistic combination of velocities using only elementary concepts from special relativity. However, some readers may find a more mathematically rigorous explanation clearer.

### v12

@@ 16.1 Manifestly Covariant Formulation
The covariant equation of motion for a relativistic charged particle, subject to the four-force $K_{\mu}=eF^{\mu\nu}dx_{\nu}/d\tau$, as given in Eq. (21), represents a relativistic extension of Newton’s second law. While the classical form $md\vec{v}/dt=\vec{f}$ remain valid in the particle’s instantaneous comoving Lorentz frame, the relativistic formulation embeds these three spatial equations into the four-dimensional Minkowski spacetime.  [^53] The notion of embedding relies on the fact that Newton’s second law can always be applied in any Lorentz frame in which the particle is instantaneously at rest [64]. That is, if a comoving Lorentz frame is defined at a given instant, the particle’s evolution can be accurately predicted in this frame over an infinitesimal time interval. Geometrically, Newton’s law holds rigorously on a hyperplane orthogonal to the particle’s world line. As the particle moves, the hyperplane—and its normal vector $u_{\mu}$—tilts accordingly along the world line. To formalize this embedding, one introduces a projection operator $\hat{P}_{\perp}$ that continuously projects vectors in Minkowski space onto hyperplanes orthogonal to the world line. This operator is given by $(\hat{P}_{\perp})_{\mu\nu}=\eta_{\mu\nu}-u_{\mu}u_{\nu}/u^{2}$ [64]. The presence of $\hat{P}_{\perp}$ highlights that only three of the four components of the covariant equation are independent.
@@ 16.2 Conventional Particle Tracking
It should be noted that usually the notation $\tau$” arose from the proper time. Now let us recall the standard concept of the object’s proper time. Let a point-like object move uniformly and rectilinearly relative to the lab frame $K$. The proper frame $K^{\prime}$ can be fixed to the moving object. The object is at rest in this frame, so that events happening with this object are registered by one clock. This clock counts the proper time at the point where the object is located. As we already know, the proper time on moving object flows slower than the time $t$. This phenomenon was called time dilation. The proper time can also be introduced for a point particle moving with acceleration. This standard interpretation of the proper time $\tau$ has nothing to do with dynamics and one may wonder where this contradiction existing between the name and the content of time in dynamic law comes from.
@@ 16.4 Standard Integration of the 4D Covariant Equation of Motion
We begin by analyzing the textbook’s treatment of the integration of the covariant equation of motion. Consider, for example, the motion of a charged particle in a given electromagnetic field. Of particular practical importance is the simplest case: a uniform electromagnetic field, meaning that the field tensor $F^{\mu\nu}$ remains constant throughout the entire space-time region of interest. Specifically, we focus on the motion of a particle in a constant, homogeneous magnetic field, represented by the tensor components $F^{\mu\nu}=B(e^{\mu}_{2}e^{\nu}_{3}-e^{\nu}_{2}e^{\mu}_{3})$ where $e^{\mu}_{2}$ and $e^{\mu}_{3}$ are orthonormal space like basis vectors $e^{2}_{2}=e^{2}_{3}=-1$, $e_{2}\cdot e_{3}=0$. In the lab frame of reference where $e^{\mu}_{0}$ is taken as the time axis, and $e^{\mu}_{2}$ and $e^{\mu}_{3}$ are space vectors the field is indeed purely magnetic, of magnitude $B$ and parallel to the $e_{1}$ axis. Let us set the initial four-velocity $u^{\mu}(0)=\gamma ce^{\mu}_{0}+\gamma ve^{\mu}_{2}$, where $v$ is the initial particle’s velocity relative to the lab observer along the axis $e_{2}$ at the instant $\tau=0$, and $\gamma=1/\sqrt{1-v^{2}/c^{2}}$. The components of the equation of motion are then $du^{(0)}/d\tau=0$, $du^{(1)}/d\tau=0$, $du^{(2)}/d\tau=-eBu^{(3)}/(mc)$, $du^{(3)}/d\tau=eBu^{(2)}/(mc)$. We seek the initial value solution to these equations as done in the existing literature  [^54] Textbook authors, such as those in [5, 64, 65], conclude from the covariant equations that a charged particle in a constant magnetic field moves in a uniform circular trajectory. This analysis does not account for relativistic kinematics, and instead, the Galilean vector law of velocity addition is applied. The non-relativistic kinematics are derived from the relation $d\tau=dt/\gamma$. It is only after making this substitution for $d\tau$ hat the authors obtain the familiar formula for the non-covariant trajectory.. A distinctive aspect of relativistic dynamics is the presence of constraints. The evolution of the particle is subject to $mdu^{\mu}/d\tau=eF^{\mu\nu}u_{\nu}$, but also to the constraint $u^{2}=c^{2}$. However, such a condition can be weakened requiring its validity at certain values of $\tau$ only, let us say initially, at $\tau=0$. Therefore, if $Y(\tau)=u^{2}-c^{2}$ vanishes initially, i.e. $Y(0)=0$, then $Y(\tau)=0$ at any $\tau$. In other words, the differential Lorentz-force equation implies the constraint $u^{2}=c^{2}$ once this is satisfied initially. Integrating with respect to the proper time we have $u^{\mu}(\tau)=\gamma e^{\mu}_{0}+\gamma v[e^{\mu}_{2}\cos(\omega\tau)+e^{\mu}_{3}\sin(\omega\tau)]$ where $\omega=eB/(mc)$. We see that $\gamma$ is constant with time, meaning that the energy of a charged particle moving in a constant magnetic field is constant. After two successive integrations we have $X^{\mu}(\tau)=X^{\mu}(0)+\gamma c\tau e^{\mu}_{0}+R[e^{\mu}_{2}\sin(\omega\tau)-e^{\mu}_{3}\cos(\omega\tau)]$ where $R=\gamma v/\omega$. This enables us to find the time dependence $0,X^{(2)}(t),X^{(3)}(t)]$ of the particle’s position since $t/\gamma=\tau$. From this solution, we conclude that the motion of a charged particle in a constant magnetic field is indeed uniform circular motion, consistent with classical expectations, yet derived within the fully covariant relativistic framework.
@@
One might expect that the particle’s trajectory in the lab frame, derived from the previous reasoning $0,X^{(2)}(t),X^{(3)}(t)]$, should be identified with $\vec{x}_{cov}(t)$. However, the trajectory $0,X^{(2)}(t),X^{(3)}(t)]$ does not include relativistic kinematics effects. We have found that the standard integration of the four-dimensional covariant equation of motion, as shown in Eq. (21), yields a particle trajectory that is mathematically identical to the one derived from Newtonian kinematics and dynamics. Specifically, the trajectory of the electron fails to incorporate relativistic effects, and instead, the Galilean law of velocity addition is applied.
@@ 16.5 Convention-Invariant Particle Tracking
16.6 Einstein’s Velocities Addition Vs. Covariant Addition
@@
This is a good point to make some general remarks about Einstein’s velocity addition. It should be noted that presented above the commonly accepted derivation of velocity composition does not follow from the composition of Lorentz boosts. In particular, second boost from $K^{\prime}$ to particle proper frame does not Lorentz boost. Einstein’s velocity addition is based on the relation $dt^{\prime}=dt/\gamma$, which implicitly assumes an absolute time convention for second boost. However, it does not account for the interplay between transverse positions $y^{\prime}$ and time $t^{\prime}$, which is inherent in Lorentz boost matrices.
In contrast, every Lorentz boost represents a transformation between reference frames within a Lorentzian framework (i.e., under Lorentz coordinatization including also transverse direction). While Lorentz boosts intertwine space and time coordinates, Einstein’s velocity addition relies on the assumption of a shared three-dimensional space across all inertial frames. The consequence of this difference is the Wigner rotation accompanying the composition of non-parallel 3D velocities.
One of the consequences of the non-commutativity of non-collinear Lorentz boosts is the absence of a shared, common notion of ordinary space. Suppose that in the Lorentz frame $K^{\prime}$, a traveler observes a particle moving along the $y^{\prime}$ axis. In other words, the particle has a velocity component $v^{\prime}_{y}$ in this frame, while the frame $K^{\prime}$ itself is moving with velocity $V$ along the $x$ axis relative to initial Lorentz frame $K$. The proper frame is rotated by an angle $\theta_{w}=v^{\prime}_{y}/(\gamma c)$ relative to frame $K$. The coordinate axes of the proper frame remain parallel to those of frame $K^{\prime}$. From the perspective of frame $K$, the coordinate axes of frame $K^{\prime}$ are also rotated by the angle $\theta_{w}$ with respect to those of frame $K$.
Let us consider the simple case where the boost velocity $V$ approaches the speed of light (i.e., the ultrarelativistic limit), while the transverse velocity $v^{\prime}_{y}$ remains small. This means we are taking the limit where $\gamma\gg 1\gg v^{\prime}_{y}/c$, treating $v^{\prime}_{y}/c$ as a small parameter and keeping terms up to second order. In this regime, the axes of the $K^{\prime}$ frame appear rotated with respect to those of $K$ by an angle $\theta_{w}=v^{\prime}_{y}/(\gamma c)$. Notably, this rotation is in the same direction as the rotation of the particle’s velocity vector in $K$. We now consider the physical consequences of this rotation. First, note that the transverse velocity component $v^{\prime}_{y}$ acquires a projection onto the $x$-axis when viewed from frame $K$. As a result, the component of the particle’s velocity along the $x$-axis in frame $K$ is slightly reduced from $V$.
@@
This result stands in contrast with the textbook prediction, which states that the total speed of the particle in frame $K$ increases from $V$ to $V+v_{y}^{2}/(2c)$. Our result contradicts this claim. The standard derivation of velocity addition neglects the mixing of transverse spatial coordinate $y^{\prime}$ with time $t^{\prime}$ in Lorentz transformations. This omission leads to an incorrect conclusion about the magnitude of the resulting velocity in initial inertial frame $K$.
It is important to highlight another effective approach to covariant particle tracking. Here, we revisit the motion of an accelerated relativistic particle from the perspective of an inertial lab observer—without changing reference frames. Specifically, we consider a particle undergoing acceleration within the lab frame, representing an example of an active boost. In this scenario, we track the system’s evolution entirely within a single (lab) reference frame.
The simplest synchronization method maintains the same set of uniformly synchronized clocks used before the acceleration, without any adjustments. This approach is based on the absolute time (or absolute simultaneity) convention, which preserves simultaneity across space. Under this convention, the standard Galilean velocity addition rule can be applied instant by instant to track the particle’s motion as it accelerates.
As discussed in Chapter 3, absolute time coordinatization can be transformed into Lorentz coordinatization. By combining Galilean transformations with appropriate changes of variables, we can derive the Lorentz transformation within the framework of absolute time. To further illustrate this idea, we now examine how distant lab clocks resynchronize during the particle’s acceleration. This analysis reveals a direct connection between the transformation of particle velocity (in Lorentz coordinates) and time dilation.
As previously emphasized, the Lorentz coordinate system is a conceptual construct. The notion of synchronized clocks in this system is purely hypothetical, used primarily for the application of Maxwell’s equations in radiation calculations.
Suppose that, before acceleration, we choose a Lorentz coordinate system in the lab frame. Immediately after the acceleration, particle velocity changes by an infinitesimal amount $d\vec{v}$ along the $y$-axis. If clock synchronization is fixed, this corresponds to adopting the absolute time convention. To maintain Lorentz coordinates in the lab frame—as discussed previously—we must perform a clock resynchronization by introducing an infinitesimal time shift. The simplest case arises when the $v_{y}$ is very small, allowing us to work up to second-order terms in $v_{y}^{2}/c^{2}$. This restriction greatly simplifies the calculations for two reasons. First, relativistic corrections due to the composition of non-collinear velocity increments only appear at order $v_{y}^{3}/c^{3}$ and can be neglected. Second, time dilation effects also enter only at higher orders.
To maintain a Lorentz coordinate system in the lab frame after the acceleration, a resynchronization of clocks is required. In this context, operations involving the rule clock structure in the lab frame can be interpreted as a change of variables governed by the transformation in Eq. 13: $y_{L}=\gamma_{y}y$, $t_{L}=(t/\gamma_{y}+\gamma_{y}yv_{y}/c^{2})$, where $\gamma_{y}=1+v_{y}^{2}/(2c^{2})$ and the particle’s displacement is given by $y=v_{y}t$.
This transformation effectively rescales the time variable—adjusting the synchronization of all clocks from $t$ to $\gamma_{y}t$, with $\gamma_{y}\simeq 1+\theta_{k}^{2}/2$. As a result, we see that the particle’s speed after acceleration is unaffected by its transverse motion. Although no second-order relativistic correction appears in the transverse ($y$-axis) velocity component, the longitudinal component is modified. Specifically, the longitudinal velocity $V$ changes to $V/\gamma_{y}$ with $v_{z}=V-v_{y}^{2}/(2c)$.
Thus, after clock resynchronization, the total electron speed in the lab frame remains VV—a result that aligns perfectly with the outcome derived earlier using a sequence of Lorentz transformations in changing frames. All pieces fit together seamlessly. We will explore the topic of relativistic velocity composition in greater detail in the next chapter
17 Mathematical Analysis of Relativistic Velocity Composition
@@
In this chapter, we shall explore the subject of relativistic velocity addition more mathematically. Textbook authors often assume that the principle of relativity requires Einstein’s velocity addition law to obey group composition. However, the non-associativity of non-collinear velocity addition is not widely recognized. Many physicists, having studied relativity from standard textbooks, might find this claim surprising, since Einstein’s velocity addition is commonly assumed to be associative. Yet, a straightforward algebraic exercise reveals its non-associative nature.
@@
The root cause of this issue lies in the presence of the Wigner rotation. In fact, the breakdown of commutativity and associativity in Einstein’s velocity addition law naturally manifests as Wigner rotation, offering deeper insight into the structure of relativistic transformations.
17.1 Parametrization of Lorentz Transformations
Einstein formally defined the special theory of relativity through two explicit postulates. Minkowski, in 1908, recognized that these postulates could be reformulated as a single geometric axiom: spacetime possesses a pseudo-Euclidean geometry. As a consequence of this spacetime geometry, Lorentz coordinate systems are related via Lorentz transformations.